OpenStudy (anonymous):

use the discriminant to find all values of b for which the equation has one real solution. x^2-bx+4=0 use the discriminant to find all values of b for which the equation has one real solution. x^2-bx+4=0 @Mathematics

OpenStudy (anonymous):

equation has one real solution only when discriminant is 0 D=b^2-4ac=b^2-4*1*4=b^2-16 b^2-16=0 (b-4)(b+4)=0 b-4=0 b=-4 or, b+4=0 b=-4 so, b=4 and b=-4